OpenStudy (anonymous):

find an equation for the line with the points 3,-7 and -6,9 in the general form Ax+By=C

OpenStudy (anonymous):

help

OpenStudy (mertsj):

Did you find the slope of the line?

OpenStudy (anonymous):

no

OpenStudy (mertsj):

That would be the first step. Do you know how to find the slope?

OpenStudy (anonymous):

yeah

OpenStudy (mertsj):

Ok. Do it and tell me what you get.

OpenStudy (anonymous):

-16/9

OpenStudy (mertsj):

That's what I got.

OpenStudy (mertsj):

Now pick one of the points. Which one do you choose?

OpenStudy (anonymous):

3,-7

OpenStudy (mertsj):

\[y-y _{1}=m(x-x _{1})\]

OpenStudy (mertsj):

Put -7 in place of y1. Put -16/9 in place of m and put 3 in place of x1

OpenStudy (mertsj):

\[y-(-7)=-\frac{16}{9}(x-3)\] \[y+7=\frac{-16}{9}x+\frac{16}{3}\] \[y=\frac{-16}{9}x-\frac{5}{3}\]

OpenStudy (mertsj):

Is that what you got?

OpenStudy (anonymous):

no, if you minus 7, u have to minus it from both sides

OpenStudy (anonymous):

that , and its not in the form Ax+By=C

OpenStudy (mertsj):

Whoops. I forgot about that part. Let's rearrange it.

OpenStudy (anonymous):

the actual answer is 16x+9y=-15, but i dont know how to get there

OpenStudy (mertsj):

\[y=\frac{-16}{9}x-\frac{5}{3}\] Multiply both sides by 9 to clear the fractions

OpenStudy (mertsj):

\[9(y)=9(\frac{-16}{9})x-9(\frac{-5}{3})\]

OpenStudy (mertsj):

\[9y=-16x+15\]

OpenStudy (mertsj):

Correction: 9y=-16x-15

OpenStudy (mertsj):

16x+9y=-15

OpenStudy (anonymous):

how did you get 5-3?

OpenStudy (anonymous):

i mean 5/3

OpenStudy (mertsj):

16/3-7

OpenStudy (mertsj):

\[y+7=\frac{-16}{9}x+\frac{16}{3}\]

OpenStudy (mertsj):

subtract 7 from both sides in that equation.

OpenStudy (mertsj):

16/3-7=-5/3

OpenStudy (anonymous):

so u use point slope form to find b, then rearange to fit Ax+By=c?

OpenStudy (mertsj):

yes. or you could use slope-intercept form if you prefer.

OpenStudy (anonymous):

how would slope intercept work

OpenStudy (mertsj):

\[y=mx+b\] \[7=\frac{-16}{9}(3)+b\]

OpenStudy (mertsj):

Make the 7 negative.

OpenStudy (mertsj):

\[-7=\frac{-16}{3}+b\]

OpenStudy (mertsj):

Add 16/3 to both sides.

OpenStudy (mertsj):

\[\frac{-5}{3}=b\]

OpenStudy (mertsj):

\[y=\frac{-16}{9}x-\frac{5}{3}\]

OpenStudy (mertsj):

do you understand?

OpenStudy (anonymous):

awsome, thank you

OpenStudy (mertsj):

yw. Now practice some more problems so you can fix it in your mind.